# Typical Positivity of Nonequilibrium Entropy Production for Pure States

### v1

@@ Abstract
Here, we clarify this dissonance by showing that “most” pure quantum states give rise to an EP “very close” to the ensemble picture (in a sense made precise below) without any of the assumptions in Refs. von Neumann (1929, 2010); Van Kampen (1954); Tasaki (2000a); Goldstein et al. (2013); Iyoda et al. (2017, 2022); Strasberg et al. (2023); Nagasawa et al. (2024); Hokkyo and Ueda (2024).
Our results rely on dynamical typicality Bartsch and Gemmer (2009); Müller et al. (2011); Reimann (2016); Balz and Reimann (2017); Reimann (2018); Dabelow and Reimann (2020); Reimann and Gemmer (2020); Dabelow and Reimann (2021); Xu et al. (2022); Teufel et al. (2023a, b, 2024) and certain continuity properties mentioned below. In particular, we will use a measure concentration result proven by Teufel, Tumulka and Vogel (Corollary 1 in Ref. Teufel et al. (2024)), which can be seen as a generalization of Levy’s lemma Milman and Schechtman (2001); Popescu et al. (2006) to non-uniform distributions.
@@
The Scrooge measure $\mu_{\rho}$ can be generated by sampling pure states $|\psi\rangle\sim\sqrt{\rho}|\phi\rangle$ with $|\phi\rangle$ a Haar random state, yielding $\rho={\langle{|\psi{\rangle\langle}\psi|}\rangle}_{\text{Scrooge}}$ (with ${\langle{\dots}\rangle}_{\text{Scrooge}}$ the average over $\mu_{\rho}$). Given $\rho$, it is the most spread out ensemble because every (complete orthogonal rank-1) projective measurement yields the same information Jozsa et al. (1994). It further arises naturally for open systems when sampling pure system-bath states in an suitable unbiased way Goldstein et al. (2006, 2016). Thus, since the Scrooge ensemble does not contain any information beyond $\rho$, it is natural to use it in the following to sample pure states. However, towards the end we will discuss other sampling schemes and we will explicitly compare it with the prominent two-point measurement scheme (TPMS) Esposito et al. (2009); Campisi et al. (2011); Deffner and Campbell (2019); Strasberg (2022); Potts (2024).
We continue by analytically showing and numerically verifying our idea for the important setup of two finite systems exchanging energy, which has direct relevance, e.g., for transport experiments with cold atoms Brantut et al. (2012, 2013); Häusler et al. (2021). Luckily, it turns out that the argument is generalizable to other setups, as we discuss afterwards. Another setup is also explicitly treated in the supplemental material Sup .
@@
Within the ensemble picture one assumes that the initial state $\rho(0)=\pi_{A}(0)\otimes\pi_{B}(0)$ is a product of two Gibbs states at (in general) different temperatures $T_{A}(0)\neq T_{B}(0)$. Note that this choice does not imply $\rho_{X}(t)=\tau_{X}(t)$ (with $X\in\{A,B\}$) at later times $t>0$ owing to the interaction $V$. Under these conditions the following inequality was proven Strasberg and Winter (2021); Strasberg et al. (2021):
@@
For the pure state version of this second law we sample $|\psi(0)\rangle$ from the Scrooge measure $\mu_{\rho(0)}$ with $\rho(0)=\pi_{A}(0)\otimes\pi_{B}(0)$. Moreover, we set $d_{X}^{-1}\equiv\|\tau_{X}(0)\|$ and argue that $d_{X}$ can be interpreted as an (often extremely large) effective dimension of $X$ as follows. First, for any density matrix $\rho$ we have $\|\rho\|=\lambda_{\text{max}}$ with $\lambda_{\text{max}}$ the largest eigenvalue of $\rho$. Next, owing to ensemble equivalence, a Gibbs state for a many-body system closely resembles a microcanonical state for which $\lambda_{\text{max}}=1/D_{\text{mic}}$ with $D_{\text{mic}}={\cal{O}}(10^{N})$ the dimension of the microcanonical energy shell, which is exponentially large in the particle number $N$. Thus, for instance for the cold atoms experiments of Refs. Brantut et al. (2012, 2013); Häusler et al. (2021) one can reasonably assume $d_{X}={\cal{O}}(10^{10000})$. Finally, quantities computed from $|\psi(t)\rangle=e^{-iHt}|\psi(0)\rangle$ get a superscript $\psi$ such as $E_{X}^{\psi}(t)$, ${\cal{S}}_{X}^{\psi}(t)$ and $\beta_{X}^{\psi}(t)$ to distinguish them from the respective quantities appearing in the ensemble picture. The stochastic EP is then defined as
@@
$\begin{split}\mu_{\rho(0)}&\left\{\big{|}E_{X}^{\psi}(t)-E_{X}(t)\big{|}>N_{X}\epsilon\right\}\\ &\leq 12\exp\left(-\frac{C\epsilon^{2}N_{X}^{2}}{\|H_{X}\|^{2}}d_{A}d_{B}\right).\end{split}$ (4)
Here, we introduced the particle number $N_{X}$ to get the right scaling: since internal energy scales extensively, it matters that we can choose an $\epsilon\ll 1$ such that the energy densities are close with high probability. This is now guaranteed: since $\|H_{X}\|\sim N_{X}$, the right hand side scales as $e^{-d_{A}d_{B}}\sim e^{-\exp(N_{A}+N_{B})}$ with respect to $N_{X}$, thus strongly suppressing fluctuations away from the ensemble averaged density $E_{X}(t)/N_{X}$ even for moderate values of $N_{X}$. Importantly, Eq. (4) holds for all times $t$ (this is dynamical typicality).
@@
$\begin{split}\mu_{\rho(0)}&\left\{\big{|}S_{X}^{\psi}(t)-S_{X}(t)\big{|}>N_{X}\epsilon\frac{c^{\prime}}{c}\frac{1}{T_{X}(t)}\right\}\\ &\leq 12\exp\left(-\frac{C\epsilon^{2}N_{X}^{2}}{\|H_{X}\|^{2}}d_{A}d_{B}\right).\end{split}$ (8)
@@
Our result closes a gap by curing deficiencies of both the ensemble picture approach and previous pure state second laws. In fact, while the ensemble picture approach proves $\Sigma\geq 0$, it does not prove that entropy gets maximized and it does not exclude arbitrary strong and frequent recurrences down to $\Sigma=0$. Conversely, while von Neumann’s $H$-theorem von Neumann (1929, 2010) proves the eventual maximization of entropy, it does not allow conclusions about transient dynamics. Our result implies that transient second-law-violating states are exceptionally rare (but in unison with time reversal symmetry not impossible), and it suggests that temporary violations of the second law, or recurrences of the EP, are unlikely unless the dynamics guides the state into atypical second-law-violating subspaces. This behaviour cannot be excluded in the ensemble picture approach because a repeated use of the Gibbs state assumption would violate unitarity of the dynamics, but—as we showed—it only matters that the pure state is sufficiently representative of a Gibbs state at later times, see also Van Kampen (1954); Strasberg et al. (2023).
Finally, the present approach could be experimentally more meaningful than the TPMS with its unrealistic perfect energy resolution. For the future it is worth to ask whether there is a universal function $f(\sigma)$ such that a fluctuation theorem ${\langle{f(\sigma)}\rangle}=1$ holds for the Scrooge ensemble. Certainly, owing to the much stronger suppression of fluctuations in $\sigma$ compared to the TPMS, the conventional choice $f(\sigma)=e^{-\sigma}$ will not suffice. Thus, besides its foundational implications, our results highlight crucial differences between different sampling procedures. This motivates the search for experimentally realistic ones and explicitly demonstrates that quantum stochastic thermodynamics is a much richer framework than its classical counterpart.
@@ Appendix A Supplemental material
A.0.1 Detailed fluctuation theorem for Clausius’ inequality

### v2

@@ Abstract
Here, we clarify this dissonance by showing that “most” pure quantum states give rise to an EP “very close” to the ensemble picture (in a sense made precise below) without any of the assumptions in Refs. von Neumann (1929, 2010); Van Kampen (1954); Tasaki (2000a); Goldstein et al. (2013); Iyoda et al. (2017, 2022); Strasberg et al. (2023); Nagasawa et al. (2024); Hokkyo and Ueda (2024). Our results rely on dynamical typicality Bartsch and Gemmer (2009); Müller et al. (2011); Monnai and Yuasa (2014); Reimann (2016); Balz and Reimann (2017); Reimann (2018); Dabelow and Reimann (2020); Reimann and Gemmer (2020); Dabelow and Reimann (2021); Xu et al. (2022); Teufel et al. (2023a, b, 2024); et al. (2024) and certain continuity properties mentioned below. In particular, we will use a measure concentration result proven by Teufel, Tumulka and Vogel (Corollary 1 in Ref. Teufel et al. (2024)), which can be seen as a generalization of Levy’s lemma Milman and Schechtman (2001); Popescu et al. (2006) to non-uniform distributions.
@@
The Scrooge measure $\mu_{\rho}$ can be generated by sampling pure states $|\psi\rangle\sim\sqrt{\rho}|\phi\rangle$ with $|\phi\rangle$ a Haar random state, yielding $\rho={\langle{|\psi{\rangle\langle}\psi|}\rangle}_{\text{Scrooge}}$ (with ${\langle{\dots}\rangle}_{\text{Scrooge}}$ the average over $\mu_{\rho}$). Given $\rho$, it is the most spread out ensemble because every (complete orthogonal rank-1) projective measurement yields the same information Jozsa et al. (1994). It further arises naturally for open systems when sampling pure system-bath states in an suitable unbiased way Goldstein et al. (2006, 2016). Thus, since the Scrooge ensemble does not contain any information beyond $\rho$, it is natural to use it in the following to sample pure states. The smallness of the bound (1) is controlled by the smallness of $\|\rho\|$, which equals the maximum eigenvalue $\lambda_{\text{max}}$ of $\rho$. Its inverse $d_{\text{eff}}\equiv 1/\|\rho\|$ can be seen as an effective dimension of the system, describing how much $\rho$ is spread out over the Hilbert space; e.g., for a maximally mixed (microcanonical) state over some (sub)space ${\cal{H}}$ of dimension $D$ one finds $d_{\text{eff}}=D$. This picture is also supported by noticing that $\|\rho\|^{2}\leq P(\rho)\leq\|\rho\|$ with $P(\rho)=\mbox{tr}\{\rho^{2}\}$ the purity of $\rho$. Thus, $\|\rho\|$ is small, or spread out over a large $d_{\text{eff}}$, if and only if $\rho$ is highly mixed.
We now continue by analytically showing and numerically verifying our idea for the important setup of two finite systems exchanging energy, which has direct relevance, e.g., for transport experiments with cold atoms Brantut et al. (2012, 2013); Häusler et al. (2021). An explicit numerical comparison with the prominent two-point measurement scheme (TPMS) Esposito et al. (2009); Campisi et al. (2011); Deffner and Campbell (2019); Strasberg (2022); Potts (2024) is also given. Afterwards, we discuss possible caveats and ways to generalize the argument to other setups and sampling schemes. Another setup is also explicitly treated in the supplemental material Sup .
@@
Within the ensemble picture one assumes that the initial state $\rho(0)=\pi_{A}(0)\otimes\pi_{B}(0)$ is a product of two Gibbs states at (in general) different temperatures $T_{A}(0)\neq T_{B}(0)$. Note that this choice does not imply $\rho_{X}(t)=\pi_{X}(t)$ (with $X\in\{A,B\}$) at later times $t>0$ owing to the interaction $V$. Under these conditions the following inequality was proven Strasberg and Winter (2021); Strasberg et al. (2021):
@@
For the pure state version of this second law we sample $|\psi(0)\rangle$ from the Scrooge measure $\mu_{\rho(0)}$ with $\rho(0)=\pi_{A}(0)\otimes\pi_{B}(0)$. Computing $d_{\text{eff}}^{X}=1/\|\pi_{X}(0)\|$ is not possible without knowing $H$, but estimates are readily obtained. For instance, equivalence of ensembles implies that $\pi_{X}(0)$ is similar to a microcanonical state, whose $d_{\text{eff}}^{X}={\cal{O}}(10^{N})$ scales exponentially with the particle number $N$. Likewise, if $X$ is composed of $N$ non-interacting (fermionic, bosonic, spin, etc.) modes $\mu$ each with purity $P(\pi_{\mu})$, we find for the total purity $P[\pi_{X}(0)]=\prod_{\mu=1}^{N}P(\pi_{\mu})$, which also has an exponential scaling with $N$. Thus, for instance for the cold atoms experiments of Refs. Brantut et al. (2012, 2013); Häusler et al. (2021) one can reasonably assume $d_{\text{eff}}^{X}={\cal{O}}(10^{10000})$. Finally, quantities computed from $|\psi(t)\rangle=e^{-iHt}|\psi(0)\rangle$ get a superscript $\psi$ such as $E_{X}^{\psi}(t)$, ${\cal{S}}_{X}^{\psi}(t)$ and $\beta_{X}^{\psi}(t)$ to distinguish them from the respective quantities appearing in the ensemble picture. The stochastic EP is then defined as
@@
$\begin{split}\mu_{\rho(0)}&\left\{\big{|}E_{X}^{\psi}(t)-E_{X}(t)\big{|}>N_{X}\epsilon\right\}\\ &\leq 12\exp\left(-\frac{C\epsilon^{2}N_{X}^{2}}{\|H_{X}\|^{2}}d_{\text{eff}}^{A}d_{\text{eff}}^{B}\right).\end{split}$ (4)
Here, we introduced the particle number $N_{X}$ to get the right scaling: since internal energy scales extensively, it matters that we can choose an $\epsilon\ll 1$ such that the energy densities are close with high probability. This is now guaranteed: since $\|H_{X}\|\sim N_{X}$, the right hand side scales as $e^{-d_{\text{eff}}^{A}d_{\text{eff}}^{B}}\sim e^{-\exp(N_{A}+N_{B})}$ with respect to $N_{X}$, thus strongly suppressing fluctuations away from the ensemble averaged density $E_{X}(t)/N_{X}$ even for moderate values of $N_{X}$. Importantly, Eq. (4) holds for all times $t$ (this is dynamical typicality).
@@
$\begin{split}\mu_{\rho(0)}&\left\{\big{|}S_{X}^{\psi}(t)-S_{X}(t)\big{|}>N_{X}\epsilon\frac{c^{\prime}}{c}\frac{1}{T_{X}(t)}\right\}\\ &\leq 12\exp\left(-\frac{C\epsilon^{2}N_{X}^{2}}{\|H_{X}\|^{2}}d_{\text{eff}}^{A}d_{\text{eff}}^{B}\right).\end{split}$ (8)
@@
Our result closes a gap by curing deficiencies of both the ensemble picture and previous pure state second laws. In fact, while the ensemble picture proves $\Sigma\geq 0$, it does not prove that entropy gets maximized and it does not exclude arbitrary strong and frequent recurrences down to $\Sigma=0$. Conversely, while von Neumann’s $H$-theorem von Neumann (1929, 2010) proves the eventual maximization of entropy, it does not allow conclusions about transient dynamics. Our result implies that transient second-law-violating states are exceptionally rare (but in unison with time reversal symmetry not impossible), and it suggests that temporary violations of the second law, or recurrences of the EP, are unlikely unless the dynamics guides the state into atypical second-law-violating subspaces. This behaviour cannot be excluded in the ensemble picture because a repeated use of the Gibbs state assumption would violate unitarity of the dynamics, but—as we showed—it only matters that the pure state is sufficiently representative of a Gibbs state at later times, see also Van Kampen (1954); Strasberg et al. (2023).
Our approach also sheds light on contemporary debates in quantum thermodynamics. While the ensemble picture seemingly provides a consistent thermodynamic framework for any bath size, our results indicate that a good bath needs a minimum size $d_{\text{eff}}$: if the bath is too small, the thermodynamics sensitively depends on inaccessible information about the precise state of the bath. Somewhat echoing this contrast, a definition of EP has been suggested that diverges if the bath is in a pure state Landi and Paternostro (2021) or that a pure state bath has zero temperature Elouard and Lombard Latune (2023). This conflicts with our approach and also with other recent findings O’Donovan et al. (2024).
Finally, the present approach is not restricted to weak coupling, can be applied to multiple conserved charges (including non-commuting ones Majidy et al. (2023)), might provide an experimentally more meaningful alternative to the TPMS with its unrealistic perfect energy resolution (for other alternatives see, e.g., Refs. Allahverdyan and Nieuwenhuizen (2005); Miller and Anders (2017); Strasberg (2019); Mohammady and Romito (2019); Micadei et al. (2020); Kerremans et al. (2022)), and therefore provides a useful tool for many current research directions.
@@ Appendix A Supplemental material
A.0.1 Integral fluctuation theorem for Clausius’ inequality